### v6

@@ Abstract
To find a suitable framework for nonlinear crystal optics(NCO), we have revisited linear crystal optics(LCO). At the methodological level, three widely used plane wave bases are compared in terms of eigenanalysis in reciprocal space and light field propagation in real space. Inspired by complex ray tracing, we expand M.V. Berry & M.R. Dennis’s 2003 uniform plane wave model to non-uniform Fourier crystal optics(FCO) and ultimately derive the explicit form of its 32 transition matrix, bridging the two major branches of crystal optics(CO) in reciprocal space, where either ray direction or spatial frequency serves as the input variable. Using this model, we create the material-matrix tetrahedral compass(M-M TC) to conduct a detailed analysis of how the four fundamental characteristics of materials (linear/circular birefringence/dichroism) influence the eigensystems of the vector electric field in two-dimensional(2D) spatial frequency domain and its distribution in three-dimensional(3D) space with a crystal-2f configuration. Along this journey, we have uncovered new territories in LCO in both real and reciprocal space, such as infinite singularities arranged in disk-, ring-, and crescent-like shapes, “L shorelines” resembling hearts, generalized haunting theorem, double conical refraction(DCR), and optical knots(OKs) it induces. We also present our model’s early applications in focal engineering and NCO. For focusing: dual-eigenmode (1) decomposition, explaining Raman spikes in conical refraction(CR); (2) interference, leading to longitudinal multifoci; (3) backward propagation, for non-Zernike aberration correction. For NCO: (1) chiral second harmonic conical refraction(SHCR); (2) linear and nonlinear spin-orbit coupling; (3) nonlinear optical(NO) orbital angular momentum(OAM) non-conservation — all three vector/scalar NCO phenomena, reproduced pixel by pixel. As the opening chapter in a trilogy, this work interweaves crystal optics, Fourier optics, and nonlinear optics, while integrating theoretical, computational, and experimental physics, advancing all six domains.
@@ 1 Introduction
Think about the ordinary(o) and exordinary(e) waves in nonlinear crystal optical(NCO) phase-matching types[6, 7, 8, 9, 10, 11, 12](e.g. type-I ooe in Fig.˜5a): what are they? Eigenmodes governed by CO[13, 4, 14, 15, 16, 17, 18, 19]!! That is to say, the entire NCO process is constantly subject to the constrains imposed by CO[13, 5, 3, 4], which is itself inherently complex[20]. More phenomenologically, for all frequency-mixing phenomena, both the pumps/fundamental waves(FWs) and the newly generated frequencies(NFs) within the crystal, must be first decomposed into the material’s intrinsic eigenmodes, then diffract independently (and anisotropically) during subsequently linearly superposition and interference, while potentially enduring absorption or gain[21, 22].
@@
In this Article, we propose a non-uniform linear Fourier crystal optical(LFCO) model, whose CO part integrates the advantages of both uniform[75, 76, 77, 78, 79, 80, 81, 1, 5, 20] and non-uniform[83, 15, 85, 14, 84, 86, 87, 111, 112] plane wave models , solving a pure bi-quadratic under FO and boundary constraints. This LFCO model extends the standard non-diagonal/non-positive-definite[113, 114, 115, 116, 117, 118, 119, 113, 120, 121, 122]/non-Hermitian[123, 124, 125, 126, 127, 128, 129, 130, 131, 132]/non-unitary[133, 134, 135, 136]/non-normal[107]/non-(mirror-[137])symmetric[5, 138, 139]/non-reciprocal[140, 141, 137] 22[142] transfer[98, 143]/transition[98, 110] matrix( evolution operator[144]) in CO to a 32 form (see Eqs.˜5 and 6), enabling explicit fully vectorial[93, 91, 145, 146, 147, 148, 149, 150, 151] computation within arbitrary materials, and at the same time, allows the LO part to directly operate in the non-uniform[83, 15, 85, 14, 84, 86, 87, 92, 84] FO framework[91, 62, 90, 93, 89].
To indirectly validate this LFCO model, and highlight its powerful role in advancing the analytic development of NCO, we offer a roadmap for expanding the LFCO model in Fig.˜2c. Following its conceptual architectures — scalar(, semi-vector,) and (full-)vector NFCO — we first demonstrate this LFCO model’s applications: in full-vector NFCO, via chiral second-harmonic conical refraction(SHCR) in Figs.˜1 and 2, and harmonic spin-orbit(S-O) angular momentum(AM) cascade in Fig.˜4 (see also Figs. SC3-SC5); in scalar NFCO, through phase-matching-controlled orbital angular momentum(OAM) conversion in Fig.˜3 (see also Figs. SB2-SB6); and in semi-vector NFCO, via full conical phase matching(FCPM) in Fig.˜5 (see also Fig. SC2).
To directly validate this LFCO model, as stated at the beginning of Section˜2.4, the procedure must begin with verifying the correctness of the eigenvalue-eigenvector computations (Fig. S8), followed by confirming the accuracy of the field distributions in real space (Fig.˜7, Fig.˜8, Figs. S9-S11, and Figs. SA2-SA10).
@@
Aided by crystal-2f system proposed from Fig. S9, we map out a second, more experimentally relevant[158, 159, 139, 81, 160] panorama in the 3D domain in Fig.˜7, freezing the AE of the light field distribution between the three vertices of the M-M TC, corresponding to the material’s birefringence(Bi), LD, and OA.
Still in real space, but with higher numerical aperture(N.A.), we offer a unified solution for the forward propagation and inverse design of focal fields in Fig.˜8, where a dual-eigenmode decomposition is proposed to explain Raman spikes[161] in conical refraction(CR) arising solely from slow modes, multifoci induced by laser processing inside materials, and aberration correction without Zernike polynomials. Along the way, we have observed two surprising new phenomena: double conical refraction(DCR) and the resulting optical field knots.
@@ 2.1 By 2025, it still remains a great challenge to reproduce chiral SHCR
The optically active( chiral[160]) SHCR[159, 165, 166, 162, 163] experiment in Fig.˜1f[162] represents an NCO experiment that, in principle, cannot be well-reproduced without a mature LCO framework[163, 164, 167, 168, 166, 169, 170, 171, 162, 172, 173]. Almost all previous NCO modeling attempts[170, 168, 168, 166, 167] have invariably failed to capture the experimental phenomena at a pixel-level resolution far beyond phenomenology. Other works refrain from modeling altogether[162, 163, 171].
@@ 2.2 The union of CO, NO, and FO: a theoretical, experimental, and computational Holy Grail
On a phenomenological level, this NCO process(i.e. SHCR) involves synchronous but distinct LCO anisotropic diffraction for both the fundamental and harmonic waves(HWs). In Fig.˜1f, the FW() conical refracts(CR) along its optic axis, while the HW() tends to double refract(DR) (yet mixed with CR, see Fig. S9 in Supplementary Note 8), forming a CRωCR-DR[134])2ω combination. In Fig.˜2, pumped along the harmonic’s axis, the configuration shifts to DRωDR-CR[175])2ω. Each Fourier component of each eigenmode, in both 2D spatial and one-dimensional(1D) angular frequency domains , exhibits wave vector double(DR)/conical(CR)/double-conical(DR-CR) or even double conical refraction(DCR) (see Fig.˜8d) at the material interfaces, superimposed on walk-off between Poynting vector and wave vector inside material.
@@
So, we ask, how, exactly, to calculate the Chiral SHCR process in Fig.˜1? Most previous efforts have ignored the foundation of NCO — LCO, and unceasingly phenomenologically to decouple LCO(anisotropic diffraction) from NCO(frequency conversion processes), both of which occur in parallel across all spatiotemporal spectral components, all eigenmodes, and all tensor & vector components. Some ask, parallel computation? Yes, but still either mathematically incorrect or numerically inefficient.
On a computational level, each of the six NCO phase-matching types() in Fig.˜1a,d (and Fig.˜2b) is, for chiral SHCR, phase-mismatched and “very likely” non-degenerate. Wherein, phase mismatch implies that none of these processes dominates the NCO in terms of energy conversion efficiency, necessitating the simultaneous consideration of all six. Non-degeneracy means that outcomes of the six (eigen)mode combinations() are usually distinct(see Fig.˜1d), requiring separate calculations for each.
@@
For KTP, one vector NFCO process takes scalar NFCO runs (Fig.˜1c). If one further scans wavelength (pulse injection, see Figs.˜4, 1, 2 and 3 and SC2), propagation distance (dynamical evolution, see Figs. S9-S11, Fig.˜8 and Fig.˜1c), pump power, beam waist (see Fig. S11c), incident angle (see Fig.˜2, Fig.˜3, and Fig. S9), N.A. (see Fig.˜8 and Fig. S11c), temperature , external magnetic field , -frame orientation (see Figs. SA4-SA5), or material coefficients (for adiabatic tuning, see Fig.˜7 and Fig. S8) — each adds a for-loop layer, and each layer may grow due to low crystallographic symmetry, or dense parameter sampling.
@@
The incoming second part of this trilogy (our scalar NFCO model) points out: with a : scale mismatch — crystal macro(10 mm[162]), lightwave meso(1 m) — Green-function formalism[262, 263] struggles transversely/in-plane(x-y), while split-step Fourier[264, 265, 266, 267, 268, 269] and pseudo-spectral method[270, 271, 272] (with Runge-Kutta scheme[273, 44, 274, 275]) strain longitudinally/out-of-plane(z), making forward computation of even one scalar NFCO hard, let alone all 42 in a full-vector NFCO. To thoroughly resolve the scalar NFCO process, we formulate the imminent Nonlinear Angular Spectrum Theory(NLAST), with its validity demonstrated in Fig.˜3 (and Figs. SB2-SB6).
All forms of N(F)CO modeling — scalar (Fig.˜3 and Figs. SB2-SB6), semi-vector (Fig.˜5 and Figs. SC2), or full-vector (Figs.˜1, 2 and 4 and Figs. SC3-SC5) — depend on a vector L(F)CO model (Figs.˜6, 7 and 8, Figs. S8-S11, and Figs. SA2-SA10) to yield the requisite eigenvectors (eigen-polarizations) and eigenvalues (eigen-wavevector-z-components for FO) , necessary for evaluating the initial complex amplitudes (i.e. modal coefficients) in Eq. (S56c) to calculate frequency-mixing dynamics, and for determining the nonlinear conversion efficiency “eigenvalue mask” sinc “eigenvector mask” in Fig.˜5.
To showcase the basics of semi-vector NFCO, Fig.˜5j reproduces the full conical phase matching(FCPM) second harmonic generation(SHG) proposed by Belyi et al.[221], illustrating the calculation of conversion in a uniaxial BBO crystal through Fig.˜5a-i, where both the first- and second-order susceptibilities are anisotropic.
In subsequent (both scalar and vector) NFCO, transverse wave vector conservation (Fig.˜5a,b) is elevated to the same fundamental level as conservation and phase continuity at the boundaries, making it a prerequisite that must take priority. The conversion efficiency, determined solely by the longitudinal phase mismatch (Fig.˜5f) and effective nonlinear coefficient (Fig.˜5i)[277, 13, 3, 4]
is subordinated to the fulfillment of transverse momentum conservation. Notably, , , and in Eq.˜1 represents the eigen-polarization states of , , and in the frame, whose combination are fixed under certain phase-matching type. For example, phase-matching type-I ooe from Fig.˜5g,h needs to compute = from Eq.˜1. Above -frame eigenvectors are almost directly provided by Berry & Dennis’s 2003 model[5] in Fig.˜10b, though a few minor adjustments in Fig.˜10c are still required.
As two masks determining the angular distribution of frequency conversion efficiency, the degree of longitudinal phase matching sinc from Fig.˜5c and the effective nonlinear coefficient from Fig.˜5i, respectively, depend on the eigenvalues and eigenvectors in linear Fourier crystal optics(LFCO), reaffirming that NCO is fundamentally grounded in LCO, and both linear(LO) and nonlinear optics(NO) should ultimately be incorporated into the framework of Fourier optics(FO).
@@ 2.3 NO’s two pillars: CO (32 transition matrix) and FO (FT pairs OTFs)
which is a transcendental equation of with being provided by various even-spectrum( biquadratic) LCO models[74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20]. Eq.˜3 can be solved through direct iteration (see Method), Newton’s iteration[241], or other methods (see Discussion and Supplementary Note 6.2), with convergence typically occurring within two iterations when anisotropy is weak and the N.A. is small.
The eigenvalues and eigenvectors (i.e., eigen-polarization states) of non-uniform LFCO can be obtained by respectively substituting obtained through Eq.˜3 into
and in Method. Then, as depicted in Fig.˜6 and Method, the eigenvalue pairs construct the propagation matrix , while the eigenvector pairs form the eigen-polarization state matrices and . A sandwich multiplication of these three matrices yields the 32 transition matrix (see Supplementary Note 5)
which structurally resembles the matrix exponential after Jordan decomposition (see Supplementary Note 1), realizing the transition from the two-component vector field on the input plane at to the three-component vector field on the output plane at through . Coupled with Fourier transform(FT) pairs (defined in Eqs. (S4, S5)) , Eq.˜5 allows us to study the evolution and 3D distribution of the vector electric field in arbitary anisotropic dielectrics in real space, as demonstrated in Fig.˜6.
Apart from CO’s in-crystal 32 transition matrix in Eq.˜5 and FO’s FT pairs in Fig.˜6, this LFCO model and the full trilogy, make broad use of self-built out-of-crystal optical transfer functions(OTFs) for lenses (Figs.˜1, 2, 4, 3, 7 and 8 and Fig. S9), objectives (Figs.˜8 and 4), half- and quarter-wave plates (Fig.˜4), q-plates (Fig.˜4), linear polarizers/analyzers (Fig. S11), etc., all within a custom FO framework.
@@
The field itself varies over space (2D reciprocal space, for eigensystems in Fig. S8; 3D real space, for wave diffraction in Fig.˜7). This work first explores the field’s adiabatic evolution(AE) (Fig. S8 and Fig.˜7) or dynamic evolution (Fig. S9 and Fig.˜8) under parameter shifts, using the proposed M-M TC( M- TC in this work), implying a “change of the change”.
Arbitrary complex dielectric tensors are typically associated with birefringent-chiral-dichroic dielectrics[153, 279, 280]. In order to relate the mathematical properties of the matrix to the optical/physical properties of the material, we propose the M-M TC in Fig.˜7 and Fig. S8.
@@
The potentially complex in Eq.˜2 is generally non-diagonalizable in 7 out of 11 cases, unless is a normal matrix[107], corresponding to the 4 possible combinations, that is, OA + Bi, LD + CD, Bi + LD, OA + CD, i.e. two endpoints of the M-M TC’s 4 edges ‘H’, ‘!H’, ‘S’, ‘!S’, respectively.
@@
Having mastered the computation of eigensystems within arbitrary materials, the next task is to batch-reproduce camera-recorded data in real space. For this purpose, both the in-crystal CO eigenmodes with their 32 transition matrix (Eqs.˜5 and 6) in reciprocal space and specific out-of-crystal FO OTFs are essential.
@@
We now revisit the real-space M-M TC in Fig.˜7, upon completing the quantitative validation in Fig. S9 around the optic axis, i.e. the diabolic point of a typical biaxial material, where the eigenvalue degeneracy does not extend to the eigenvectors.
@@
Perform a counterclockwise scan along the three edges of the M-M TC. First, increasing OA from 0 along the ‘H’ edge of the M-M TC yields the AE diagram of the optical rotation field from vertice ‘Bi’ to ‘OA’ in Fig.˜7e (pointed by the light blue arrows), ultimately converging to Bloembergen et al.’s experiment: chiral CR[159] through -HIO3. Next, while maintaining OA, as LD increases along the ‘Im’ edge from 0 (denoted by the light green arrows), Brenier’s experimental results on chirality versus dichroism[160] of acentric chiral Nd3+-doped BZBO are obtained, with a clear criterion for OA LD. That is, when pump’s polarization is aligned parallel to the polarization of Voigt wave while the analyzer is set perpendicular, central extinction is observed in the far field. Following this, by canceling OA and keeping only LD, the angular absorption distribution of laser crystal KGd(WO4)2 is obtained[139, 81]. Then, by gradually reducing LD along the ‘S’ edge (antiparallel to the light red arrows), the Pancharatnam phenomenon[158] emerges.
Finally, as LD is reduced to 0, only Bi remains, including uniaxiality (Fig. S10b and Fig. S11), biaxiality (Fig. S10a and Fig. S11b1), and their hyperbolic counterparts (Fig. S10a,b). As an effect caused by the off-diagonal elements of the 22 eigen-polarization matrix and an embodiment of conservation law, the global grasp and detailed calculation of SOI play a significant role in not only LCO[215, 216, 217, 219, 221, 223, 222, 224, 225, 226] (see Fig.˜6 and Figs. S9, S11 and SA3-SA5) but also NCO[247, 281, 6, 255] (see Fig.˜4).

### v7

@@ Abstract
To find a suitable framework for nonlinear crystal optics(NCO), we have revisited linear crystal optics(LCO). At the methodological level, three widely used plane wave bases are compared in terms of eigenanalysis in reciprocal space and light field propagation in real space. Inspired by complex ray tracing, we expand M.V. Berry and M.R. Dennis’s 2003 uniform plane wave model to non-uniform Fourier crystal optics and ultimately derive the explicit form of its 32 transition matrix, bridging the two major branches of crystal optics in reciprocal space, where either ray direction or spatial frequency serves as the input variable. Using this model, we create the material-matrix tetrahedral compass to conduct a detailed analysis of how the four fundamental characteristics of materials (linear/circular birefringence/dichroism) influence the eigensystems of the vector electric field in two-dimensional spatial frequency domain and its distribution in three-dimensional space with a crystal-2f configuration. Along this journey, we have uncovered new territories in LCO in both real and reciprocal space, such as infinite singularities arranged in disk-, ring-, and crescent-like shapes, “L shorelines” resembling hearts, generalized haunting theorem, double conical refraction, and optical knots it induces. We also present our model’s early applications in focal engineering and NCO. As the opening chapter in a trilogy, this work interweaves crystal optics, Fourier optics, and nonlinear optics, while integrating theoretical, computational, and experimental physics, advancing all six domains.
@@ 1 Introduction
Think about the ordinary(o) and exordinary(e) waves in nonlinear crystal optical(NCO) phase-matching types[6, 7, 8, 9, 10, 11, 12](e.g. type-I ooe in Fig. 5a): what are they? Eigenmodes governed by CO[13, 4, 14, 15, 16, 17, 18, 19]!! That is to say, the entire NCO process is constantly subject to the constrains imposed by CO[13, 5, 3, 4], which is itself inherently complex[20]. More phenomenologically, for all frequency-mixing phenomena, both the pumps/fundamental waves(FWs) and the newly generated frequencies(NFs) within the crystal, must be first decomposed into the material’s intrinsic eigenmodes, then diffract independently (and anisotropically) during subsequently linearly superposition and interference, while potentially enduring absorption or gain[21, 22].
@@
In this Article, we propose a non-uniform linear Fourier crystal optical(LFCO) model, whose CO part integrates the advantages of both uniform[75, 76, 77, 78, 79, 80, 81, 1, 5, 20] and non-uniform[83, 15, 85, 14, 84, 86, 87, 111, 112] plane wave models , solving a pure bi-quadratic under FO and boundary constraints. This LFCO model extends the standard non-diagonal/non-positive-definite[113, 114, 115, 116, 117, 118, 119, 113, 120, 121, 122]/non-Hermitian[123, 124, 125, 126, 127, 128, 129, 130, 131, 132]/non-unitary[133, 134, 135, 136]/non-normal[107]/non-(mirror-[137])symmetric[5, 138, 139]/non-reciprocal[140, 141, 137] 22[142] transfer[98, 143]/transition[98, 110] matrix( evolution operator[144]) in CO to a 32 form (see Eqs. 5 and 6), enabling explicit fully vectorial[93, 91, 145, 146, 147, 148, 149, 150, 151] computation within arbitrary materials, and at the same time, allows the LO part to directly operate in the non-uniform[83, 15, 85, 14, 84, 86, 87, 92, 84] FO framework[91, 62, 90, 93, 89].
To indirectly validate this LFCO model, and highlight its powerful role in advancing the analytic development of NCO, we offer a roadmap for expanding the LFCO model in Fig. 2c. Following its conceptual architectures — scalar(, semi-vector,) and (full-)vector NFCO — we first demonstrate this LFCO model’s applications: in full-vector NFCO, via chiral second-harmonic conical refraction(SHCR) in Figs. 1 and 2, and harmonic spin-orbit(S-O) angular momentum(AM) cascade in Fig. 4 (see also Figs. SC3-SC5); in scalar NFCO, through phase-matching-controlled orbital angular momentum(OAM) conversion in Fig. 3 (see also Figs. SB2-SB6); and in semi-vector NFCO, via full conical phase matching(FCPM) in Fig. 5 (see also Fig. SC2).
To directly validate this LFCO model, as stated at the beginning of Section 2.4, the procedure must begin with verifying the correctness of the eigenvalue-eigenvector computations (Fig. S8), followed by confirming the accuracy of the field distributions in real space (Fig. 7, Fig. 8, Figs. S9-S11, and Figs. SA2-SA10).
@@
Aided by crystal-2f system proposed from Fig. S9, we map out a second, more experimentally relevant[158, 159, 139, 81, 160] panorama in the 3D domain in Fig. 7, freezing the AE of the light field distribution between the three vertices of the M-M TC, corresponding to the material’s birefringence(Bi), LD, and OA.
Still in real space, but with higher numerical aperture(N.A.), we offer a unified solution for the forward propagation and inverse design of focal fields in Fig. 8, where a dual-eigenmode decomposition is proposed to explain Raman spikes[161] in conical refraction(CR) arising solely from slow modes, multifoci induced by laser processing inside materials, and aberration correction without Zernike polynomials. Along the way, we have observed two surprising new phenomena: double conical refraction(DCR) and the resulting optical field knots.
@@ 2.1 By 2025, it still remains a great challenge to reproduce chiral SHCR
The optically active( chiral[160]) SHCR[159, 162, 163, 164, 165] experiment in Fig. 1f[164] represents an NCO experiment that, in principle, cannot be well-reproduced without a mature LCO framework[165, 166, 167, 168, 163, 169, 170, 171, 164, 172, 173]. Almost all previous NCO modeling attempts[170, 168, 168, 163, 167] have invariably failed to capture the experimental phenomena at a pixel-level resolution far beyond phenomenology. Other works refrain from modeling altogether[164, 165, 171].
@@ 2.2 The union of CO, NO, and FO: a theoretical, experimental, and computational Holy Grail
On a phenomenological level, this NCO process(i.e. SHCR) involves synchronous but distinct LCO anisotropic diffraction for both the fundamental and harmonic waves(HWs). In Fig. 1f, the FW() conical refracts(CR) along its optic axis, while the HW() tends to double refract(DR) (yet mixed with CR, see Fig. S9 in Supplementary Note 8), forming a CRωCR-DR[134])2ω combination. In Fig. 2, pumped along the harmonic’s axis, the configuration shifts to DRωDR-CR[175])2ω. Each Fourier component of each eigenmode, in both 2D spatial and one-dimensional(1D) angular frequency domains , exhibits wave vector double(DR)/conical(CR)/double-conical(DR-CR) or even double conical refraction(DCR) (see Fig. 8d) at the material interfaces, superimposed on walk-off between Poynting vector and wave vector inside material.
@@
So, we ask, how, exactly, to calculate the Chiral SHCR process in Fig. 1? Most previous efforts have ignored the foundation of NCO — LCO, and unceasingly phenomenologically to decouple LCO(anisotropic diffraction) from NCO(frequency conversion processes), both of which occur in parallel across all spatiotemporal spectral components, all eigenmodes, and all tensor & vector components. Some ask, parallel computation? Yes, but still either mathematically incorrect or numerically inefficient.
On a computational level, each of the six NCO phase-matching types() in Fig. 1a,d (and Fig. 2b) is, for chiral SHCR, phase-mismatched and “very likely” non-degenerate. Wherein, phase mismatch implies that none of these processes dominates the NCO in terms of energy conversion efficiency, necessitating the simultaneous consideration of all six. Non-degeneracy means that outcomes of the six (eigen)mode combinations() are usually distinct(see Fig. 1d), requiring separate calculations for each.
@@
For KTP, one vector NFCO process takes scalar NFCO runs (Fig. 1c). If one further scans wavelength (pulse injection, see Figs. 4, 1, 2 and 3 and SC2), propagation distance (dynamical evolution, see Figs. S9-S11, Fig. 8 and Fig. 1c), pump power, beam waist (see Fig. S11c), incident angle (see Fig. 2, Fig. 3, and Fig. S9), N.A. (see Fig. 8 and Fig. S11c), temperature , external magnetic field , -frame orientation (see Figs. SA4-SA5), or material coefficients (for adiabatic tuning, see Fig. 7 and Fig. S8) — each adds a for-loop layer, and each layer may grow due to low crystallographic symmetry, or dense parameter sampling.
@@
The incoming second part of this trilogy (our scalar NFCO model) points out: with a : scale mismatch — crystal macro(10 mm[164]), lightwave meso(1 m) — Green-function formalism[264, 265] struggles transversely/in-plane(x-y), while split-step Fourier[266, 267, 268, 269, 270, 271] and pseudo-spectral method[272, 273, 274] (with Runge-Kutta scheme[275, 44, 276, 253]) strain longitudinally/out-of-plane(z), making forward computation of even one scalar NFCO hard, let alone all 42 in a full-vector NFCO. To thoroughly resolve the scalar NFCO process, we formulate the imminent Nonlinear Angular Spectrum Theory(NLAST), with its validity demonstrated in Fig. 3 (and Figs. SB2-SB6).
All forms of N(F)CO modeling — scalar (Fig. 3 and Figs. SB2-SB6), semi-vector (Fig. 5 and Figs. SC2), or full-vector (Figs. 1, 2 and 4 and Figs. SC3-SC5) — depend on a vector L(F)CO model (Figs. 6, 7 and 8, Figs. S8-S11, and Figs. SA2-SA10) to yield the requisite eigenvectors (eigen-polarizations) and eigenvalues (eigen-wavevector-z-components for FO) , necessary for evaluating the initial complex amplitudes (i.e. modal coefficients) in Eq. (S56c) to calculate frequency-mixing dynamics, and for determining the nonlinear conversion efficiency “eigenvalue mask” sinc “eigenvector mask” in Fig. 5.
To showcase the basics of semi-vector NFCO, Fig. 5j reproduces the full conical phase matching(FCPM) second harmonic generation(SHG) proposed by Belyi et al.[221], illustrating the calculation of conversion in a uniaxial BBO crystal through Fig. 5a-i, where both the first- and second-order susceptibilities are anisotropic.
In subsequent (both scalar and vector) NFCO, transverse wave vector conservation (Fig. 5a,b) is elevated to the same fundamental level as conservation and phase continuity at the boundaries, making it a prerequisite that must take priority. The conversion efficiency, determined solely by the longitudinal phase mismatch (Fig. 5f) and effective nonlinear coefficient (Fig. 5i)[277, 13, 3, 4]
is subordinated to the fulfillment of transverse momentum conservation. Notably, , , and in Eq. 1 represents the eigen-polarization states of , , and in the frame, whose combination are fixed under certain phase-matching type. For example, phase-matching type-I ooe from Fig. 5g,h needs to compute = from Eq. 1. Above -frame eigenvectors are almost directly provided by Berry & Dennis’s 2003 model[5] in Fig. 10b, though a few minor adjustments in Fig. 10c are still required.
As two masks determining the angular distribution of frequency conversion efficiency, the degree of longitudinal phase matching sinc from Fig. 5c and the effective nonlinear coefficient from Fig. 5i, respectively, depend on the eigenvalues and eigenvectors in linear Fourier crystal optics(LFCO), reaffirming that NCO is fundamentally grounded in LCO, and both linear(LO) and nonlinear optics(NO) should ultimately be incorporated into the framework of Fourier optics(FO).
@@ 2.3 NO’s two pillars: CO (32 transition matrix) and FO (FT pairs OTFs)
which is a transcendental equation of with being provided by various even-spectrum( biquadratic) LCO models[74, 75, 76, 77, 78, 79, 80, 81, 1, 5, 20]. Eq. 3 can be solved through direct iteration (see Method), Newton’s iteration[241], or other methods (see Discussion and Supplementary Note 6.2), with convergence typically occurring within two iterations when anisotropy is weak and the N.A. is small.
The eigenvalues and eigenvectors (i.e., eigen-polarization states) of non-uniform LFCO can be obtained by respectively substituting obtained through Eq. 3 into
and in Method. Then, as depicted in Fig. 6 and Method, the eigenvalue pairs construct the propagation matrix , while the eigenvector pairs form the eigen-polarization state matrices and . A sandwich multiplication of these three matrices yields the 32 transition matrix (see Supplementary Note 5)
which structurally resembles the matrix exponential after Jordan decomposition (see Supplementary Note 1), realizing the transition from the two-component vector field on the input plane at to the three-component vector field on the output plane at through . Coupled with Fourier transform(FT) pairs (defined in Eqs. (S4, S5)) , Eq. 5 allows us to study the evolution and 3D distribution of the vector electric field in arbitary anisotropic dielectrics in real space, as demonstrated in Fig. 6.
Apart from CO’s in-crystal 32 transition matrix in Eq. 5 and FO’s FT pairs in Fig. 6, this LFCO model and the full trilogy, make broad use of self-built out-of-crystal optical transfer functions(OTFs) for lenses (Figs. 1, 2, 4, 3, 7 and 8 and Fig. S9), objectives (Figs. 8 and 4), half- and quarter-wave plates (Fig. 4), q-plates (Fig. 4), linear polarizers/analyzers (Fig. S11), etc., all within a custom FO framework.
@@
The field itself varies over space (2D reciprocal space, for eigensystems in Fig. S8; 3D real space, for wave diffraction in Fig. 7). This work first explores the field’s adiabatic evolution(AE) (Fig. S8 and Fig. 7) or dynamic evolution (Fig. S9 and Fig. 8) under parameter shifts, using the proposed M-M TC( M- TC in this work), implying a “change of the change”.
Arbitrary complex dielectric tensors are typically associated with birefringent-chiral-dichroic dielectrics[153, 279, 280]. In order to relate the mathematical properties of the matrix to the optical/physical properties of the material, we propose the M-M TC in Fig. 7 and Fig. S8.
@@
The potentially complex in Eq. 2 is generally non-diagonalizable in 7 out of 11 cases, unless is a normal matrix[107], corresponding to the 4 possible combinations, that is, OA + Bi, LD + CD, Bi + LD, OA + CD, i.e. two endpoints of the M-M TC’s 4 edges ‘H’, ‘!H’, ‘S’, ‘!S’, respectively.
@@
Having mastered the computation of eigensystems within arbitrary materials, the next task is to batch-reproduce camera-recorded data in real space. For this purpose, both the in-crystal CO eigenmodes with their 32 transition matrix (Eqs. 5 and 6) in reciprocal space and specific out-of-crystal FO OTFs are essential.
@@
We now revisit the real-space M-M TC in Fig. 7, upon completing the quantitative validation in Fig. S9 around the optic axis, i.e. the diabolic point of a typical biaxial material, where the eigenvalue degeneracy does not extend to the eigenvectors.
@@
Perform a counterclockwise scan along the three edges of the M-M TC. First, increasing OA from 0 along the ‘H’ edge of the M-M TC yields the AE diagram of the optical rotation field from vertice ‘Bi’ to ‘OA’ in Fig. 7e (pointed by the light blue arrows), ultimately converging to Bloembergen et al.’s experiment: chiral CR[159] through -HIO3. Next, while maintaining OA, as LD increases along the ‘Im’ edge from 0 (denoted by the light green arrows), Brenier’s experimental results on chirality versus dichroism[160] of acentric chiral Nd3+-doped BZBO are obtained, with a clear criterion for OA LD. That is, when pump’s polarization is aligned parallel to the polarization of Voigt wave while the analyzer is set perpendicular, central extinction is observed in the far field. Following this, by canceling OA and keeping only LD, the angular absorption distribution of laser crystal KGd(WO4)2 is obtained[139, 81]. Then, by gradually reducing LD along the ‘S’ edge (antiparallel to the light red arrows), the Pancharatnam phenomenon[158] emerges.
Finally, as LD is reduced to 0, only Bi remains, including uniaxiality (Fig. S10b and Fig. S11), biaxiality (Fig. S10a and Fig. S11b1), and their hyperbolic counterparts (Fig. S10a,b). As an effect caused by the off-diagonal elements of the 22 eigen-polarization matrix and an embodiment of conservation law, the global grasp and detailed calculation of SOI play a significant role in not only LCO[215, 216, 217, 219, 221, 223, 222, 224, 225, 226] (see Fig. 6 and Figs. S9, S11 and SA3-SA5) but also NCO[247, 281, 6, 256] (see Fig. 4).